The temperature in your body is 750F. If the answer is 23.90C in Celsius scale and 296.9K in Kelvin scale, how will you get these answers?

1
Expert's answer
2021-11-21T17:25:17-0500

To convert temperature in Fahrenheit to Celsius, use this equation:


"(T_F-32)·\frac59=(75-32)·\frac59=23.9°\text C."

To convert temperature in Fahrenheit to Kelvin, use this equation:


"(T_F-32)·\frac59+273=(75-32)·\frac59+273=\\=296.9\text K."